OpenStudy (anonymous):

What is the ratio 4 to 60 expressed as a fraction in simplest form

OpenStudy (leader):

\[\frac{ 4 }{ 60}=\frac{ 2 }{ 30 }=\frac{ 1 }{ 15 }\]

OpenStudy (midhun.madhu1987):

What is common to both 4 and 60?

OpenStudy (leader):

remember a:b is the same thing as a/b